please help me with this problem I can't get past it! It needs to be solved using elimination and graphing method
x+y=152
x=y+28

Answer by Alan3354(69443)

x+y=152
x=y+28
=======
x+y = 152
x-y = 28
---------- Add
2x = 180
x = 90
---
x - y = 28
y = 62
===========
To graph:
They're both straight lines so you need 2 points.
Pick 2 values for y, find x, or vice versa
Plot the 2 points, draw a line thru them.
---------
The intersection of the 2 lines is the solution of the system.